Write the equation of the line passing through (-3,-5) and (3,0).

Thanks

Answer by checkley71(8403)

(-3,-5)(3,0) slope (0+5)/(3+3)=5/6 & the y intercept is -3*5/6 or -2.5 thus the equation for this line is Y=5/6X-2.5